OpenStudy (anonymous):

my answer is 4√3x 3√8x - √2x

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

I did it on mathway. No, according to mathway it 5√2x.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I also need √27x + √12x - √3x

OpenStudy (anonymous):

No problem. I dont really know how to solve it, but you should double check your work to see what you had wrong. Go here it will solve for what you need. www.mathway.com or try www.wolframalpha.com

OpenStudy (anonymous):

3√8x - √2x = 3*√4*√2 * √x - √2* √x

OpenStudy (anonymous):

And you work it out from there

OpenStudy (anonymous):

\[4\sqrt{3x}\]

OpenStudy (anonymous):

Ill try wolfram just in case mathway gave an error.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wolfram gives the same so I dont think your answer is write. Ill check the other also.

OpenStudy (anonymous):

fml what was the answer before?

OpenStudy (anonymous):

Your answer for the first one is actually correct for the second one you gave

OpenStudy (anonymous):

4 sqrt 3x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

dope. thanks.

OpenStudy (anonymous):

NP!

OpenStudy (anonymous):

love you

OpenStudy (anonymous):

:) DITTO!

OpenStudy (anonymous):

Isn't it supposed to be 5 √2x?